Dirk's age is now 5 more than twice Lebron's age. In 8 years, Dirk's age then will be 5 times Lebron's age 4 years ago. How old is Lebron now?
Answer by indra89811(24)

Let Lebron' current age be L
Dirk's present age = 2L +5
2L+5+8=5(L-4)
2L+13=5L-20
3L=33
L=11
Ans: 11 years
--------------------Peace out, hun :) ---------------------------